# POD-Attention: Unlocking Full Prefill-Decode Overlap for Faster LLM Inference

### v1

@@ 3.2. Analysis of Readily Available Methods
For kernel-parallel execution, shown as FA_Streams in Figure 6, we run FA’s prefill and decode kernel on two different CUDA streams. For warp-parallel execution (FA_HFuse), we fuse FA’s kernels using the toolchain provided by (HFuse_CGO_2022, ). Figure 6 compares their performance against serial execution of FA’s prefill and decode attention kernels (FA_Serial). Our experiment shows the per-layer attention computation time of Yi-6B for 32 chunks of a 16K prompt (chunk size 512), each co-scheduled with decodes of 16K context length each.
@@
Note that if the number of CTAs in a kernel is not divisible by the number of GPU SMs, some of the SMs in the last wave of scheduling can remain idle — a phenomenon known as wave quantization (cusync2024, ; streamk2024, ). In the worst case, a marginal increase in work can double the latency of a kernel due to wave quantization. Therefore, to fully understand the benefit of concurrent execution, we evaluate performance with and without wave quantization. Each decode request uses 4 CTAs in our experiment (one CTA per KV head). Hence a decode batch size of 54 uses 216 CTAs having no wave quantization on our NVIDIA A100 GPU (108 SMs). In contrast, a batch size of 55 uses 220 CTAs leaving 4 quantized CTAs.
@@ 3.3. Analysis of Other Methods
For complex kernels, such as attention, efficiently implementing fine-grained fusion schemes is non-trivial and prone to errors. Therefore, we analyze the performance of CTA-parallel and intra-thread fusion methods with a simple micro-benchmark consisting of a compute-bound kernel that repeatedly multiplies array elements with a scalar, and a memory-bound kernel that repeatedly adds three arrays. Each thread executes a barrier after each operation. We vary the number of compute iterations to evaluate performance under varying compositions of compute-bound and memory-bound operations. Figure 7 shows the runtime of different fusion methods applied on these two functions. At 100 compute iterations, both operations consume equal time when executed serially. To the left of this point, memory bound is more dominant. To the right, it is compute bound. Figure 7 also shows the runtime achievable with an ideal oracle (i.e., perfect overlap).
CTA-parallel cannot guarantee SM-level co-location of compute-bound and memory-bound operations and hence provides only marginal average improvement of 8% over serial execution. Intra-thread fusion outperforms both serial and CTA-parallel execution, on average by 19% and 10%. However, the benefit of intra-thread fusion is limited due to sync barriers that hinder concurrent execution.
Figure 7. Fine-grained fusion versus serial computation.
@@ 4. POD-Attention
We introduce POD-Attention — a single GPU kernel that efficiently computes both prefill and decode attention. Our primary goal is to ensure that each GPU SM computes both operations simultaneously while minimizing resource contention between them. We build our kernel atop FA v2.6.1 (flashattention2, ).
@@ 4.1. SM-aware CTA Scheduling
SM-aware CTA scheduling co-locates prefill and decode CTAs through “runtime operation binding”. Here, a CTA decides whether to perform prefill or decode at runtime, after checking: 1) which SM it got launched on (smcentric:ICS:2015, ), and 2) what other CTAs running on the same SM are doing. This allows the kernel to remain completely agnostic to how the hardware scheduler assigns SMs to CTAs.
@@
Figure 9 shows a simple code snippet of SM-aware CTA scheduling. When the hardware scheduler schedules a new CTA on an SM, a leader thread of the CTA (e.g., thread 0) reads the SMID hardware counter (PTX_SMID, ) that contains the unique ID of the SM it was launched on (lines 2 - 3). The thread then performs an atomic add operation on the SM counter to obtain a ticket (line 6). This ticket informs the thread as to which operation it should perform i.e., prefill or decode (lines 7 - 8), depending on the scheduling policy. The thread also increments the CTA counter for the operation (line 10). If this exceeds the maximum CTAs for that operation, it switches operations (line 12 - 18). Finally, it writes this information to shared memory so that the other threads in the CTA can begin execution accordingly (lines 20 - 30). We examined two scheduling policies: 50:50 and proportional. In the 50:50 policy, subsequent CTAs on an SM alternate between prefill and decode. In contrast, the proportional policy (line 5) allocates CTAs based on the ratio of prefill and decode CTAs in the current batch.
@@ 4.2.1. Tile Sizes
Data tiling is necessary to make effective use of tensor cores, which provide $\sim$8×8$\times$8 × higher throughput than their CUDA core counterpart (tensorcore, ). Tiling also helps improve shared memory usage. However, the benefit of tiling is not uniform across operations. Decode operates on a single token per request, having a tile length of one across the query sequence length (QSL) dimension. In Group Query Attention (ainslie2023gqa, ), this length increases to the ratio between query and KV heads, typically 2 – 8. Due to this small dimension length, data reuse is insignificant, and performance is limited by memory bandwidth.
@@
FlashAttention uses tile lengths of 64 – 128 for the QSL dimension. The side-effect of using such large tile sizes is that decodes end up zero padded, causing redundant compute (flashdecoding++, ). For example, 10(a) shows that compute utilization of the decode attention kernel is proportional to tile sizes, reaching up to 70% at QSL tile dimension of 128, compared to 10% with tile dimension of 16. However, note that decode attention is memory bound and hence, the primary objective of a decode kernel is to try and saturate memory bandwidth. 10(b) shows that even at a relatively large QSL tile dimension of 64, the decode kernel is able to maximize memory bandwidth utilization. Hence, for a decode-only attention kernel, there is little incentive to reducing tile sizes further.
@@ 4.2.4. Limiting Prefill Splits
FlashAttention parallelizes computation across the query heads and QSL tile dimension. FlashDecoding (flashdecoding, ), designed for decode which has a QSL of one, further splits the computation across the K/V dimension when there is not enough parallelism to fill the SMs of the GPU. The side-effect of this approach is that different CTAs fetch the same query tensor from memory independently of each other, proportional to the number of splits. Consequently, splitting the computation increases memory bandwidth utilization. While splitting along the key/value dimension is not required for prefills when the input contains enough tokens, chunked-prefills limit the number of tokens processed per-iteration by design (to minimize TBT). Therefore, FlashAttention also uses the FlashDecoding technique to accelerate the chunked-prefill attention computation. This scheme works well for a prefill-only kernel as increased parallelism can easily offset the cost of extra memory reads.
@@ 5. Evaluation
Models and environment: We evaluate POD-Attention with Yi-6B (4 KV heads (yi-6b-200k-hf, )), Llama-2-7B (32 KV heads (llama-2-7b-hf, )) and Llama-3-8B (8 KV heads (llama-8b-hf, )), deploying Yi-6B on one A100 GPU, and others on two A100 GPUs with tensor parallelism. Each model has 32 query heads. Each GPU has 80GB memory.
Workloads and metrics: We evaluate both offline and online inference scenarios. For offline inference, we report the number of requests processed per minute. For online inference, we report TTFT, TBT and request execution latency on two workloads consisting of 2K requests each, and context length ranging from 4K to 32K tokens per-request. One of the workloads is an internal enterprise workload (mean context length of 10.5K tokens, per-request prefill to decode token ratio i.e., P:D in the range of 0 – 40) and the other is based on arXiv-Summarization (arxiv, ) (mean context length of 9.5K tokens, P:D ratio of 0-50). On average, the number of decode tokens in arXiv workload is 42% higher (470) than the internal workload (331).
Serving system baselines: Our experiments use Sarathi-Serve (sarathiserve:github, ) as the serving framework, which is built atop vLLM (vLLM:github, ) . We evaluate two baselines: 1) the original vLLM scheduler (vllmsosp, ) that runs prefills and decodes in separate batches, prioritizing prefills over decodes and 2) Sarathi-Serve (sarathiserve2024, ). Both baselines use FlashAttention kernels (v2.6.1) for attention computation. We integrate POD-Attention into Sarathi-Serve to evaluate the benefits of our optimizations. For simplicity, we refer to Sarathi-Serve without and with POD-Attention as Sarathi and Sarathi+POD.
5.1. Accelerating Attention with POD-Attention
@@
In addition to FlashAttention kernels, we also compare the runtime of FlashInfer (FI) kernels (flashinfer, ) in two configurations: FI_Serial and FI_Batched. FI_Batched computes prefill and decode attention using the prefill kernel of FlashInfer. We compare against FI_Batched for two reasons: 1) this strategy is the easiest way to compute prefill and decode attention together, and 2) some systems prefer this method e.g., Sarathi used FI_Batched in its default attention back-end (sarathi:unifiedattentionkernel, ), and a similar feature is requested in vLLM (vllm:unifyattentionkernel, ). However, we show that this strategy is inefficient e.g., when FI_Batched uses a prefill-optimized kernel, it leads to redundant compute in decode computation due to use of larger tile sizes (§4.2.1). This redundant computation interferes with co-running prefill. Similarly, interference would occur on memory-bandwidth if FI_Batched uses a decode-optimized kernel.
Figure 11 shows the relative speedup for different mechanisms compared to FA_Serial. FA_Streams provides limited speedup as it cannot guarantee SM-level overlap of operations. In rare cases, we find that the overhead of stream synchronization can also negate its benefits. FI_Serial is mostly on par with FA_Serial but FlashInfer’s prefill kernel is somewhat slower than that of FlashAttention (vattention2024, ). FI_Batched improves performance at low context lengths, but degrades at higher lengths by up to 40% due to redundant computation for decodes. FA_HFuse is the strongest baseline as it guarantees operation overlap, improving median performance by 11%. However, FA_HFuse is susceptible to the straggler effect due to which it is slower by up to 16% in some cases compared to FA_Serial. The straggler effect can also be seen in Figure 6 towards the later chunks where prefill becomes more dominant, making it hard to achieve perfect utilization.
@@
POD-Attention reaches a peak speedup of 75%, and a mean of 28% — higher than all alternatives. We found that in 18% of cases, it also reaches within 10% of the theoretical peak speedup, signifying near-perfect overlap. Furthermore, unlike other alternatives, POD-Attention never under-performs serial execution. These results underline the importance of a specialized attention kernel for hybrid batching based LLM inference.
@@ 5.2. Evaluating Throughput in Offline Inference
For evaluating offline inference scenarios, we run long context requests of 16K tokens each. We use chunk size 512 for Yi-6B, and 1K for both Llama-2-7B and Llama-3-8B, chosen in a way that chunking a prompt does not reduce the performance of linear operations (as recommended by Sarathi (sarathi2023, ; sarathiserve2024, )). We run 1K total requests for Yi-6B, and 2K requests each for Llama-2-7B and Llama-3-8B such that the total runtime of a single configuration is about one hour. The number of output tokens per-request is set to 2K for Yi-6B, 1K for Llama-3-8B and 256 for Llama-2-7B; we study the effect of varying prefill to decode token ratio (P:D ratio) in §5.4.4.
Figure 12 shows that Sarathi+POD delivers the best throughput: $22\%$, $20\%$ and $19\%$ higher than Sarathi, and $27\%$, $13\%$ and $12\%$ higher than vLLM, for the three models. It is worth highlighting that chunked-prefills and hybrid batching involves a tradeoff. Chunking a prompt increases attention computation time due to repeated KV cache loads: computing attention of a prefill chunk requires reading KV cache of all prior chunks (distserve2024, ). At the same time, fusing decode tokens with prefills helps execute linear operations more efficiently: model weights need not be read separately for prefills and decodes. Therefore, the relative performance of vLLM and Sarathi can vary depending on workload, model configuration and chunk size. In our experiments, Sarathi improves throughput slightly over vLLM for Yi-6B but under-performs it for Llama-2-7B and Llama-3-8B. Sarathi+POD fuses prefills and decodes in all operations to improve GPU resource utilization, thereby outperforms both baselines.
@@ 5.3. Evaluating Latency in Online Inference
We evaluate Llama-3-8B on the internal and arXiv-based workloads near the serving capacity of the system: the maximum load a system can handle while avoiding high queuing delays (sarathiserve2024, ). We evaluate 2048 requests in each workload by varying the input load based on Poisson distribution. For Sarathi and Sarathi+POD, we use chunk size of 1024 for the arXiv-based workload, and 1536 for the internal workload which is more prefill-heavy. We discuss performance on important LLM-specific latency metrics of TTFT, TBT, and end-to-end request execution latency.
Note that there is an inherent trade-off between these metrics (sarathiserve2024, ) and optimizing for one metric can severely compromise the others. For example, as will see below, vLLM prioritizes prefills and thus achieves low TTFT but sacrifices TBT, resulting in 95+of user requests experiencing one or more stalls during decode generation. On the other hand, Sarathi reduces the stalls to a small % of user requests but significantly increases TTFT compared to vLLM.

### v2

@@ 3.2. Analysis of Readily Available Methods
For kernel-parallel execution, shown as FA_Streams in Figure 6, we run FA’s prefill and decode kernel on two different CUDA streams. For warp-parallel execution (FA_HFuse), we fuse FA’s kernels using the toolchain provided by (Li et al., 2022). Figure 6 compares their performance against serial execution of FA’s prefill and decode attention kernels (FA_Serial). Our experiment shows the per-layer attention computation time of Yi-6B for 32 chunks of a 16K prompt (chunk size 512), each co-scheduled with decodes of 16K context length each.
@@
Note that if the number of CTAs in a kernel is not divisible by the number of GPU SMs, some of the SMs in the last wave of scheduling can remain idle — a phenomenon known as wave quantization (Jangda et al., 2024; Osama et al., 2023). In the worst case, a marginal increase in work can double the latency of a kernel due to wave quantization. Therefore, to fully understand the benefit of concurrent execution, we evaluate performance with and without wave quantization. Each decode request uses 4 CTAs in our experiment (one CTA per KV head). Hence a decode batch size of 54 uses 216 CTAs having no wave quantization on our NVIDIA A100 GPU (108 SMs). In contrast, a batch size of 55 uses 220 CTAs leaving 4 quantized CTAs.
@@
Figure 7. Fine-grained fusion versus serial computation.
@@ 3.3. Analysis of Other Methods
For complex kernels, such as attention, efficiently implementing fine-grained fusion schemes is non-trivial and prone to errors. Therefore, we analyze the performance of other fusion methods with a simple micro-benchmark consisting of a compute-bound kernel that repeatedly multiplies array elements with a scalar, and a memory-bound kernel that repeatedly adds three arrays. Each thread executes a barrier after each operation. We vary the number of compute iterations to evaluate performance under varying compositions of compute-bound and memory-bound operations. Figure 7 shows the runtime of different fusion methods applied on these two functions. At 100 compute iterations, both operations consume equal time when executed serially. To the left of this point, memory bound is more dominant. To the right, it is compute bound. Figure 7 also shows the runtime achievable with an ideal oracle (i.e., perfect overlap).
CTA and kernel-parallel cannot guarantee SM-level co-location of compute-bound and memory-bound operations and hence provides only marginal average improvement of 3% and 7% over serial execution. Intra-thread fusion outperforms both serial and CTA-parallel execution, on average by 13%. However, the benefit of intra-thread fusion is limited due to sync barriers that hinder concurrent execution.
@@ 4. POD-Attention
We introduce POD-Attention — a single GPU kernel that efficiently computes both prefill and decode attention. Our primary goal is to ensure that each GPU SM computes both operations simultaneously while minimizing resource contention between them. We build our kernel atop FA v2.6.1 (Dao, 2024).
@@ 4.1. SM-aware CTA Scheduling
SM-aware CTA scheduling co-locates prefill and decode CTAs through “runtime operation binding”. Here, a CTA decides whether to perform prefill or decode at runtime, after checking: 1) which SM it got launched on (Wu et al., 2015), and 2) what other CTAs running on the same SM are doing. This allows the kernel to remain completely agnostic to how the hardware scheduler assigns SMs to CTAs.
@@
Figure 9 shows a simple code snippet of SM-aware CTA scheduling. When the hardware scheduler schedules a new CTA on an SM, a leader thread of the CTA (e.g., thread 0) reads the SMID hardware counter (PTX, 2024b) that contains the unique ID of the SM it was launched on (lines 2 - 3). The thread then performs an atomic add operation on the SM counter to obtain a ticket (line 6). This ticket informs the thread as to which operation it should perform i.e., prefill or decode (lines 7 - 8), depending on the scheduling policy. The thread also increments the CTA counter for the operation (line 10). If this exceeds the maximum CTAs for that operation, it switches operations (line 12 - 18). Finally, it writes this information to shared memory so that the other threads in the CTA can begin execution accordingly (lines 20 - 30). We examined two scheduling policies: 50:50 and proportional. In the 50:50 policy, subsequent CTAs on an SM alternate between prefill and decode. In contrast, the proportional policy (line 5) allocates CTAs based on the ratio of prefill and decode CTAs in the current batch.
@@ 4.2.1. Tile Sizes
Data tiling is necessary to make effective use of tensor cores, which provide $\sim$8×8$\times$8 × higher throughput than their CUDA core counterpart (Appleyard and Yokim, 2017). Tiling also helps improve shared memory usage. However, the benefit of tiling is not uniform across operations. Decode operates on a single token per request, having a tile length of one across the query sequence length (QSL) dimension. In Group Query Attention (Ainslie et al., 2023), this length increases to the ratio between query and KV heads, typically 2 – 8. Due to this small dimension length, data reuse is insignificant, and performance is limited by memory bandwidth.
@@
FlashAttention uses tile lengths of 64 – 128 for the QSL dimension. The side-effect of using such large tile sizes is that decodes end up zero padded, causing redundant compute (Hong et al., 2024). For example, 10(a) shows that compute utilization of the decode attention kernel is proportional to tile sizes, reaching up to 70% at QSL tile dimension of 128, compared to 10% with tile dimension of 16. However, note that decode attention is memory bound and hence, the primary objective of a decode kernel is to try and saturate memory bandwidth. 10(b) shows that even at a relatively large QSL tile dimension of 64, the decode kernel is able to maximize memory bandwidth utilization. Hence, for a decode-only attention kernel, there is little incentive to reducing tile sizes further.
@@ 4.2.4. Limiting Prefill Splits
FlashAttention parallelizes computation across the query heads and QSL tile dimension. FlashDecoding (Dao et al., 2023), designed for decode which has a QSL of one, further splits the computation across the K/V dimension when there is not enough parallelism to fill the SMs of the GPU. The side-effect of this approach is that different CTAs fetch the same query tensor from memory independently of each other, proportional to the number of splits. Consequently, splitting the computation increases memory bandwidth utilization. While splitting along the key/value dimension is not required for prefills when the input contains enough tokens, chunked-prefills limit the number of tokens processed per-iteration by design (to minimize TBT). Therefore, FlashAttention also uses the FlashDecoding technique to accelerate the chunked-prefill attention computation. This scheme works well for a prefill-only kernel as increased parallelism can easily offset the cost of extra memory reads.
@@ 4.3. Implementing CTA-parallel Fusion
4.4. Discussion on Alternative Implementations
Concurrent execution is a well studied topic in GPU literature (Krieder et al., 2014; Yeh et al., 2017; Liang et al., 2015; Wang et al., 2016), and our high-level goal of overlapping prefill and decode attention computation can be achieved in multiple ways. One noteworthy strategy is based on persistent threads (Gupta et al., 2012; Zhao et al., 2023; Pai et al., 2013): in this method, one launches a pre-determined number of CTAs (enough to perfectly fill all the SMs). Persistent threads of these CTAs pull the right type of work as necessary (e.g., prefill or decode tiles). We find that this strategy also alleviates the straggler problem. However, SM-aware scheduling is still needed to decide what work (prefill or decode) to run on which persistent CTA, critical to guaranteeing operation co-location within an SM. Upon integrating it with SM-aware scheduling, we find that this strategy performs on par with our CTA-parallel fusion.
NVIDIA also provides MPS (multi-process service) (nvi, 2024b) and MIG (multi-instance GPUs) (nvi, 2024a) features to run different applications in parallel on the same GPU. However, because hybrid batching combines prefill and decode operations within a single process by design, MPS and MIG are inapplicable to our use case.
@@ 5. Evaluation
Model GPU ##\## Q Heads ##\## KV Heads ##\## Layers
Yi-6B 1 A100 32 4 32
Llama-2-7B 2 A100s 32 32 32
Llama-3-8B 2 A100s 32 8 32
Table 4. Models and hardware used for evaluation.
@@
Models and environment: We evaluate POD-Attention with Yi-6B (4 KV heads (yi-, 2024)), Llama-2-7B (32 KV heads (lla, 2024a)) and Llama-3-8B (8 KV heads (lla, 2024b)), deploying Yi-6B on one A100 GPU, and others on two A100 GPUs with tensor parallelism (Table 4). Each model has 32 query heads. Each GPU has 80GB HBM memory.
Workloads and metrics: We evaluate both offline and online inference scenarios. For offline inference, we report the number of requests processed per minute. For online inference, we report TTFT, TBT and request execution latency on two workloads consisting of 2K requests each, and context length ranging from 4K to 32K tokens per-request. One of the workloads is an internal enterprise workload (mean context length of 10.5K tokens, per-request prefill to decode token ratio i.e., P:D in the range of 0 – 40) and the other is based on arXiv-Summarization (arx, 2024) (mean context length of 9.5K tokens, P:D ratio of 0-50). On average, the number of decode tokens in arXiv workload is 42% higher (470) than the internal workload (331).
Serving system baselines: Our experiments use Sarathi-Serve (sar, 2024b) as the serving framework, which is built atop vLLM (vLL, 2024) . We evaluate two baselines: 1) the original vLLM scheduler (Kwon et al., 2023) that runs prefills and decodes in separate batches, prioritizing prefills over decodes and 2) Sarathi-Serve (Agrawal et al., 2024c). Both baselines use FlashAttention kernels (v2.6.1) for attention computation. We integrate POD-Attention into Sarathi-Serve to evaluate the benefits of our optimizations. For simplicity, we refer to Sarathi-Serve without and with POD-Attention as Sarathi and Sarathi+POD.
5.1. Evaluating Attention Computation
@@
In addition to FlashAttention kernels, we also compare the runtime of FlashInfer (FI) v0.2.0 kernels (Ye et al., 2025) in two configurations: FI_Serial and FI_Batched. FI_Batched computes prefill and decode attention using the prefill kernel of FlashInfer. We compare against FI_Batched for two reasons: 1) this strategy is the easiest way to compute prefill and decode attention together, and 2) some systems prefer this method e.g., Sarathi used FI_Batched in its default attention back-end (sar, 2024a), and a similar feature is requested in vLLM (vll, 2024b). However, we show that this strategy is inefficient e.g., when FI_Batched uses a prefill-optimized kernel, it leads to redundant compute in decode computation due to use of larger tile sizes (§4.2.1). This redundant computation interferes with co-running prefill. Similar interference occurs on memory-bandwidth if FI_Batched uses a decode-optimized kernel.
Figure 11 shows the relative speedup for different mechanisms compared to FA_Serial. FA_Streams provides limited speedup as it cannot guarantee SM-level overlap of operations. In rare cases, we find that the overhead of stream synchronization can also negate its benefits. FI_Serial has better optimized decode kernels giving it a modest improvement over FA_Serial, but it does not overlap the operations. FI_Batched improves performance at low context lengths, but degrades at higher lengths by up to 40% due to redundant computation for decodes. FA_HFuse is the strongest baseline as it guarantees operation overlap, improving median performance by 11%. However, FA_HFuse is susceptible to the straggler effect due to which it is slower by up to 13% compared to FA_Serial. The straggler effect can also be seen in Figure 6 towards the later chunks where prefill is more dominant, making it hard to achieve perfect utilization.
@@
POD-Attention reaches a peak speedup of 59%, and a mean of 28% — higher than all alternatives. We found that in 25% of cases, it also reaches within 10% of the theoretical peak speedup, signifying near-perfect overlap. Furthermore, unlike other alternatives, POD-Attention never under-performs serial execution. These results underline the importance of a specialized attention kernel for hybrid-batching-based LLM inference.
Additionally, we profiled the energy consumption of the attention kernels and observed that POD-Attention reduces energy consumption by up to 35% over FA_Serial (mean 20.5%). These savings are largely proportional to the reduction in runtime, showing that prefill-decode overlap not only improves performance but also reduces energy consumption.
@@ 5.2. Evaluating Throughput in Offline Inference
For evaluating offline inference scenarios, we run long context requests of 16K tokens each. We use chunk size 512 for Yi-6B, and 1K for both Llama-2-7B and Llama-3-8B, chosen in a way that chunking a prompt does not reduce the performance of linear operations (as recommended by Sarathi (Agrawal et al., 2023, 2024c)). We run 1K total requests for Yi-6B, and 2K requests each for Llama-2-7B and Llama-3-8B such that the total runtime of a single configuration is about one hour. The number of output tokens per-request is set to 2K for Yi-6B, 1K for Llama-3-8B and 256 for Llama-2-7B; we study the effect of varying prefill to decode token ratio (P:D ratio) in §5.4.4.
Figure 12 shows that Sarathi+POD delivers the best throughput: $22\%$, $20\%$ and $19\%$ higher than Sarathi, and $27\%$, $13\%$ and $12\%$ higher than vLLM, for the three models. It is worth highlighting that chunked-prefills and hybrid batching involves a tradeoff. Chunking a prompt increases attention computation time due to repeated KV cache loads: computing attention of a prefill chunk requires reading KV cache of all prior chunks (Zhong et al., 2024). At the same time, fusing decode tokens with prefills helps execute linear operations more efficiently: model weights need not be read separately for prefills and decodes. Therefore, the relative performance of vLLM and Sarathi can vary depending on workload, model configuration and chunk size. In our experiments, Sarathi improves throughput slightly over vLLM for Yi-6B but under-performs it for Llama-2-7B and Llama-3-8B. Sarathi+POD fuses prefills and decodes in all operations to improve GPU resource utilization, thereby outperforms both baselines.
@@ 5.3. Evaluating Latency in Online Inference
We evaluate Llama-3-8B on the internal and arXiv-based workloads near the serving capacity of the system: the maximum load a system can handle while avoiding high queuing delays (Agrawal et al., 2024c). We evaluate 2048 requests in each workload by varying the input load based on Poisson distribution. For Sarathi and Sarathi+POD, we use chunk size of 1024 for the arXiv-based workload, and 1536 for the internal workload which is more prefill-heavy. We discuss performance on important LLM-specific latency metrics of TTFT, TBT, and end-to-end request execution latency.
Note that there is an inherent trade-off between these metrics (Agrawal et al., 2024c) and optimizing for one metric can severely compromise the others. For example, as will see below, vLLM prioritizes prefills and thus achieves low TTFT but sacrifices TBT, resulting in 95+of user requests experiencing one or more stalls during decode generation. On the other hand, Sarathi reduces the stalls to a small % of user requests but significantly increases TTFT compared to vLLM.